OpenStudy (anonymous):

Please HELP! Romeo and Juliet saw each other when they were 270 feet apart. Romeo began running towards Juliet at a rate of 16ft/s. At the same moment, Juliet began running towards Romeo at a rate of 14ft/s. A. How many seconds after they started running will they meet? B. How far will Romeo have run then?

OpenStudy (anonymous):

A.9 seconds B.144ft

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

how is it 9?

OpenStudy (anonymous):

270/ 16+14

OpenStudy (anonymous):

t=d/r They both run for the same length of time. Solve the following for d, Romeo's distance:\[\frac{d}{16}=\frac{270-d}{14}\]d = 144 feet

OpenStudy (anonymous):

it is the distance divided by how much they run a second altogether(30 feet)

OpenStudy (anonymous):

to calculate how long romeo runs, you multiply the distance he runs per second by how many seconds he needs to run that is solved in part (A.)

OpenStudy (anonymous):

\[\frac{\text{length}}{\frac{\text{length}}{\text{time}}}\text{=}\text{time}\]

OpenStudy (anonymous):

B. How far will Romeo have run then? r t = d, t = d/r t = 144/16 = 9 seconds

OpenStudy (anonymous):

Mistake. Replace B. ....... with A. How many seconds after they started running will they meet? 9 seconds.

OpenStudy (anonymous):

Mistake. Replace B. ....... with A. How many seconds after they started running will they meet? 9 seconds.